give the correct number of significant figures and units to the problem below. (3.4 m + 2.78 m)/565 s=
Answer #1

= (3.4 + 2.78)/ 565

= (6.18) / 565

= 10.9*10^-3 m/s

==========

do comment in case any doubt, will reply for sure.. Goodluck